The following is the information from national income account from the hypothetical.
GNP= 2400
GI=400
NI=150
Consumption=1500
G.P.G.S=480
NY=1925
WS=1460
PI+RI of P=160
Dividend=50
GBS=15
SIC=190
N.I.I=60
G.B.T.P=260
PT and N.T.P=300
Find the following
1.NNP
2.net export
3.indirect taxes
4.coperate profit
5. Taxes-transfers
6. Personal income
7. Disposable personal income
8. Personal saving
1
Expert's answer
2021-07-08T15:19:45-0400

1. NNP

NDP = GNP - Depreciation

Net Private Investment= Gross Private Investment – Depreciation.

"150=480- Depreciation."

"Depreciation. =480-150=330\\

NNP = GNP - Depreciation = 2400-330\\=2070"


2. Net exports

NX = GNP- Consumption - Gross investment - Government purchases

"2400-1500-400-480=20"


3:indirect taxes

indirect taxes=sales taxes-revenue

"300-190=110"


4.coperate profit

coperate profit=revenue -expenses

"2400-(160+300+1460)=480"


5.Government taxes minus transfers

T = Government expenditure + Budget Surplus

"= 1400-15=415"


6. Personal income

Personal income=GNP-retained profits + transfer payments

"2400-(260+1925)=60"


7. Disposable personal income

Disposable personal income = GNP - Government taxes minus transfers

"2400-415=1945"


8. Personal Saving

Personal Savings = Disposable personal income - Consumption

"1945-1500=485"